Please help me solve this equation by substitution.
-5x + 3y = -5, -7x - 5y = -7
I keep getting a crazy fractional answer, and I don't think it's correct! Thank you!

Found 2 solutions by macston, lwsshak3:
Answer by macston(5194)

.
-5x+3y=-5 Add 5x to each side.(Equation 1)
3y=5x-5 Divide each side by 3.
y=(5x-5)/3 Use this to substitute for y in Equation 2
.
-7x-5y=-7 Substitute for y (Equation 2)
-7x-5((5x-5)/3)=-7
-7x+(-25x+25)/3=-7 Multiply each side by 3.
-21x-25x+25=-21 Subtract 25 from each side.
-46x=-46 Divide each side by -46.
x=1
.
-5x+3y=-5 Put in value for x.
-5(1)+3y=-5
-5+3y=-5 Add 5 to each side.
3y=0 Divide each side by 3
y=0
.
CHECK:
-7x-5y=-7 Put in values.
-7(1)-5(0)=-7
-7-0=-7
-7=-7

Answer by lwsshak3(11628)

Please help me solve this equation by substitution.
-5x + 3y = -5, -7x - 5y = -7
3y=-5+5x
y=-5/3+5x/3
..
-7x - 5y = -7
sub y from 1st eq
-7x-5(-5/3+5x/3)=-7
-7x+25/3-25x/3=-7
lcd:3
-21x+25-25x=-21
46x=46x
x=1
y=-5/3+5x/3
3y=-5+5x=-5+5=0
3y=0
y=0
point of intersection: (1, 0)